State of Nature
Decision Alternative Strong Demand S1 Weak Demand S2
Small complex, d1 7 5
Medium complex, d2 14 6
Large complex, d3 20 -8

Suppose PDC is optimistic about the potential for the luxury high-rise condominium complex and that this optimism leads to an initial subjective probability assessment of 0.8 that demand will be strong (S1) and a corresponding probability of 0.2 that demand will be weak (S2). Assume the decision alternative to build the large condominium complex was found to be optimal using the expected value approach. Also, a sensitivity analysis was conducted for the payoffs associated with this decision alternative. It was found that the large complex remained optimal as long as the payoff for the strong demand was greater than or equal to $17.5 million and as long as the payoff for the weak demand was greater than or equal to -$18 million.

Consider the medium complex decision. How much could the payoff under strong demand increase and still keep decision alternative d3 the optimal solution? If required, round your answer to two decimal places.

The payoff for the medium complex under strong demand remains less than or equal to $   million, the large complex remains the best decision.

Consider the small complex decision. How much could the payoff under strong demand increase and still keep decision alternative d3 the optimal solution? If required, round your answer to two decimal places.

The payoff for the small complex under strong demand remains less than or equal to $   million, the large complex remains the best decision.
Answer #1

1) Consider the medium complex decision. How much could the payoff under strong demand increase and still keep decision alternative d3 the optimal solution? If required, round your answer to two decimal places.

ANSWER:

Expected Value of Large complex, d3 = 20*0.8-8*0.2 = 14.4

Therefore, payoff under strong demand for Medium complex can be as much such that its EMV is at most equal to 14.4

Therefore, maximum payoff under strong demand for Medium complex = (14.4 - 6*0.2)/0.8 = $ 16.5 m

The payoff for the medium complex under strong demand remains less than or equal to $__16.50__million, the large complex remains the best decision.

2) Consider the small complex decision. How much could the payoff under strong demand increase and still keep decision alternative d3 the optimal solution? If required, round your answer to two decimal places.

ANSWER:

Payoff under strong demand for Small complex can be as much such that its EMV is at most equal to 14.4

Therefore, maximum payoff under strong demand for Small complex = (14.4 - 5*0.2)/0.8 = $ 16.75 m

The payoff for the small complex under strong demand remains less than or equal to $__16.75__ million, the large complex remains the best decision.